OpenStudy (dmezzullo):

What is the slope of the line 2x - 5y = -10? negative five halves five halves two fifths negative two fifths none of the above

OpenStudy (anonymous):

m = -a /b = -2/-5 = 2/5

OpenStudy (anonymous):

negative two fifths

OpenStudy (anonymous):

negative 2/5ths (:

OpenStudy (dmezzullo):

tyvm

OpenStudy (anonymous):

no prob.. always here you know(:

OpenStudy (anonymous):

Not really. First do this: 2x + 5y = -3 Subtract 2x from each side: 2x - 2x + 5y = -3 -2x 5y = -2x -3 Now what do you get when you divide both sides bvy 5?